# Score-based Neural Ordinary Differential Equations for Computing Mean Field Control Problems

### v1

@@ Abstract
Classical neural ordinary differential equations (ODEs) are powerful tools for approximating the log-density functions in high-dimensional spaces along trajectories, where neural networks parameterize the velocity fields. This paper proposes a system of neural differential equations representing first- and second-order score functions along trajectories based on deep neural networks. We reformulate the mean field control (MFC) problem with individual noises into an unconstrained optimization problem framed by the proposed neural ODE system. Additionally, we introduce a novel regularization term to enforce characteristics of viscous Hamilton–Jacobi–Bellman (HJB) equations to be satisfied based on the evolution of the second-order score function. Examples include regularized Wasserstein proximal operators (RWPOs), probability flow matching of Fokker–Planck (FP) equations, and linear quadratic (LQ) MFC problems, which demonstrate the effectiveness and accuracy of the proposed method.
@@ 1 Introduction
Score functions have been widely used in modern machine learning algorithms, particularly generative models through time-reversible diffusion (Song et al., 2021). The score function can be viewed as a deterministic representation of diffusion in stochastic trajectories (Carrillo et al., 2019). In this representation, one reformulates the Brownian motion by the gradient of the logarithm of the density function, after which deterministic trajectories involving score functions can approximate the probability density function. These properties have inspired algorithms for simulating stochastic trajectories or sampling problems that converge to target distributions (Wang et al., 2022; Lu et al., 2024). Typical applications include modeling the time evolution of probability densities for stochastic dynamics and solving control problems constrained by such dynamics.
While score functions provide powerful tools for modeling stochastic trajectories, their computations are often inefficient, especially in high-dimensional spaces. Classical methods, such as kernel density estimation (KDE) (Chen, 2017), tend to perform poorly in such settings due to the curse of dimensionality (Terrell & Scott, 1992).
@@
Related work. The normalizing flow has emerged as a powerful technique for solving inference problems, allowing for the construction of complex probability distributions in a tractable manner. This approach was popularized by works such as Rezende & Mohamed (2015); Papamakarios et al. (2021). The advent of Neural ODEs (Chen et al., 2018) has attracted considerable attention in the community, opening up a new paradigm for continuous-time deep learning models. Following this breakthrough, numerous extensions have been proposed to enhance the expressiveness and flexibility of these models, such as the augmented Neural ODEs by Dupont et al. (2019).
@@
Regarding score-based models, Song et al. (2021) studies the score-based reversible time diffusion models in generative modeling. In this case, the score function often comes from the OU process. One needs to solve the score-matching problem to learn the time-dependent score function. In addition, the first-order score dynamic from neural ODEs has been introduced by Boffi & Vanden-Eijnden (2023) to work on the simulation of the Fokker-Planck (FP) equation. Rather than using score dynamics directly to compute the probability flow ODEs associated with FP equations, they employed score-matching methods to achieve efficient computations.
@@
The study of MFC problems has become crucial in the last decade Cardaliaguet et al. (2019); Bensoussan et al. (2013); Fornasier & Solombrino (2014). MFC studies strategic decision-making in large populations where individual players interact through specific mean-field quantities. This control formulation is also useful in generative models Zhang & Katsoulakis (2023). For example, neural network-based methods have been employed to solve MFC problems. Hu & Lauriere (2023) provided a comprehensive review of neural network approaches for MFC. For first-order MFC problems, Ruthotto et al. (2020a); Huang et al. (2023) developed Neural ODE-based methods. In parallel, Ruthotto et al. (2020b) introduced numerical algorithms based on the characteristic lines of the Hamilton-Jacobi-Bellman (HJB) equation. For second-order MFC problems, neural networks have also been employed. Lin et al. (2021) utilized generative adversarial networks (GANs) to approximate minimization systems for MFC problems, where one neural network models the population density, and the other parameterizes the value function. Reisinger et al. (2024) proposed a PDE-based iterative algorithm to tackle MFC problems. Along this line, MFC problems are generalizations of dynamical optimal transport problems, known as Benamou-Brenier’s formulas Benamou & Brenier (2000a). A famous example is the Jordan-Kinderlehrer-Otto (JKO) scheme (Jordan et al., 1998), a variational time discretization in approximating the gradient drift FP equation. The one-step iteration of the JKO scheme can be viewed as an MFC problem. In machine learning computations, the neural JKO scheme (Xu et al., 2023; Vidal et al., 2023) was introduced to compute the FP equation, and Lee et al. (2024) extended this to approximate general nonlinear gradient flows through a generalized deep neural JKO scheme. This scheme approximates a deterministic MFC problem in each time interval. Other works that use machine learning methods to solve MFC problems include Dayanikli et al. (2023); Zhou et al. (2024); Dayanikli et al. (2023). Different from previous results, our work designs first- and second-order score dynamics to compute second-order MFC problems.
@@ 2 Score-based neural ODE in continuous time
Let $z_{t}\in\mathbb{R}^{d}$ be a flow of random variable for $t\geq 0$ with initialization $z_{0}=x$, which is obtained from a pushforward map $z_{t}=T(t,x)$ in Lagrangian coordinates. For simplicity, we denote $T_{t}:=T(t,\cdot)$, so that $T_{0}$ is an identity map. Throughout this paper, we assume that $T(\cdot,\cdot)$ is smooth and $T_{t}$ is invertible for all $t\geq 0$. We denote the probability distribution of $z_{t}$ by $\rho_{t}$ or $\rho(t,\cdot)$, then $\rho_{t}=T_{t\#}\rho_{0}$, where # is the pushforward operator for distributions. By the change of variable formula, $\rho$ satisfies the well known Monge–Ampère (MA) equation (Gutiérrez & Brezis, 2001)
@@
We leave the proofs for this equation and all propositions afterwards in Appendix A.
@@
$\partial_{t}H_{t}$ $=-\sum_{i=1}^{d}s_{it}\nabla_{z}^{2}f_{i}(t,z_{t})-\nabla_{z}^{2}(\nabla_{z}\cdot f(t,z_{t}))-H_{t}\nabla_{z}f(t,z_{t})-\nabla_{z}f(t,z_{t})^{\top}H_{t}\,,$ (3d)
@@
Example: Centered Gaussian distributions. To illustrate the ODE dynamics equation 3, we consider a concrete example where the random variable follows a centered Gaussian distribution. Let the pushforward map be linear in $x$, i.e. $T(t,x)=T(t)x$, where $T(t):\mathbb{R}_{+}\to\mathbb{R}^{d\times d}$ is a time-dependent matrix. If we define $A(t):\mathbb{R}_{+}\to\mathbb{R}^{d\times d}$ by $\partial_{t}T(t)=A(t)T(t)$, then the vector field of the state dynamics is $f(t,x)=A(t)x$. Now, let the initial state $z_{0}$ follow a centered Gaussian distribution $N(0,\Sigma(0))$ with covariance matrix $\Sigma(0)$. Under this linear map $T(t)$, the state $z_{t}$ remains a Gaussian distribution, with a time-dependent covariance matrix $\Sigma(t)$. In this case, $z_{t}\sim N(0,\Sigma(t))$, and the time evolution of the covariance matrix follows the matrix ODE (with the derivation in Appendix A.1)
@@
In this case, $\nabla_{x}f(t,x)=A(t)$ and $\nabla_{x}\cdot f(t,x)=\operatorname{Tr}(A(t))$, and all higher-order spatial derivatives of $f$ vanish. Therefore, the ODE system equation 3 simplifies to
@@ 3.1 A generalization of normalizing flow
Chen et al. (2018) interpreted the deep residual neural network as an ODE with each layer representing one step of the forward Euler scheme. Similar ideas have also been proposed by Haber & Ruthotto (2017). In this section, we extend these concepts and demonstrate how the first- and second-order score functions can be computed efficiently using the proposed neural ODE system.
We partition the time interval $[0,t_{\text{end}}]$ into $N_{t}$ subintervals with the length $\Delta t=t_{\text{end}}/N_{t}$ and denote the time stamps by $t_{j}=j\,\Delta t$. We apply the forward Euler scheme to discretize the ODE system equation 3. We parametrize the vector field $f(t,z;\theta)$ as a neural network with parameter $\theta$. Here, $\theta=[w_{0},w_{1},w_{2},b_{1},b_{2}]\in\mathbb{R}^{k+k\times d+d\times k+k+d}$ and
@@
where $\sigma\colon\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is a vector function with elementwise activation functions. Here $k$ is the width of the network. This is the typical structure of a neural network with one hidden layer. Then, we discretize the state dynamic $z_{t}$ through
@@ 3.2 Different algorithms for computing the score function
The expression equation 3c or equation 12 provides a way to compute the score function. In this section, we compare several algorithms for computing the score function, and demonstrate the potential efficiency of using formulas equation 3c or equation 12.
One method is to derive the score function from the pushforward map $T(t,x)$ and the MA equation directly. To be more specific, we apply the operator $\nabla_{x}$ and the logarithm on both sides of the MA equation equation 1. We obtain
@@
As an alternative, we can compute the score $s_{t}$ through the normalizing flow $l_{t}$, i.e., differentiating $l_{t}=\log\rho(t,z_{t})$ w.r.t $z_{t}$. We assume that the width of the neural network equation 8 is $k=\mathcal{O}(d)$. The score function can be reformulated as
@@
where $\dfrac{\partial z_{t}}{\partial z_{0}}$ and $\dfrac{\partial l_{t}}{\partial z_{0}}$ are obtained from auto-differentiations of deep neural network functions in equation 15. Using the chain rule, the total computational computational cost for computing one trajectory of the score function is still $\mathcal{O}(N_{t}\,d^{3})$. Both equation 17 and equation 18 require computing the inverse Jacobian of $T$ or solving related linear systems, resulting in a cubic cost in the spatial dimension. As mentioned by Chen et al. (2018) (section 4), this is the bottleneck of these methods. We leave more detailed discussions of both methods equation 17 and equation 18, and two other methods, to Appendix D.
In contrast, the high-order normalizing flow equation 12 only requires discretizing the ODE dynamics equation 3a and equation 3c, with a total cost of $\mathcal{O}(N_{t}\,d^{2})$ for computing one score trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$. As is shown in table 1, our formulation is more efficient compared with equation 17 and equation 18, especially in high dimensions. Additionally, we are able to compute the second-order score function through 13.
pushforward map equation 17 normalizing flow equation 18 high-order normalizing flow equation 12 (ours)
$\mathcal{O}(N_{t}\,d^{3})$ $\mathcal{O}(N_{t}\,d^{3})$ $\mathcal{O}(N_{t}\,d^{2})$
@@
Example: Gaussian distribution. Our formulation for the second-order score function also has advantages in the example of Gaussian distributions, where $H_{t}=\Sigma(t)^{-1}$ is exactly the inverse of the covariance matrix. Traditionally, estimating $\Sigma(t)^{-1}$ involves several steps. One needs to sample sufficiently many points $z_{0}^{(n)}$, simulate the state dynamic to obtain $z_{t}^{(n)}$, estimate the covariance matrix from these samples, and finally compute the inverse of this estimation. To achieve an error of $\varepsilon$, at least $\mathcal{O}(\varepsilon^{-2})$ samples are required for accurate covariance estimation. Compared to our second-order score dynamic equation 3d, we only need to pick $\Delta t=\mathcal{O}(\varepsilon)$ and compute $H_{t}$ through equation 13, with a total cost of $\mathcal{O}(\varepsilon^{-1})$. This results in a significantly more efficient algorithm.
@@ 4 Solving second-order MFC problems
In this section, we apply the score-based normalizing flow to solve the second-order MFC problem. In particular, we demonstrate that the MFC problem can be represented by the optimal control problem of the neural dynamical system in Proposition 1. For readers interested in the connection between MFC problems and generative AI, we refer to the work of Zhang & Katsoulakis (2023). In the context of machine learning, Neklyudov et al. (2023) also studies similar problems, in which they compute Wasserstein Lagrangian flows. In this work, we can deal with the diffusion term in MFC problems using proposed neural normalization flows.
@@ 4.1 Formulation of MFC problem
$\inf_{v}\int_{0}^{t_{\text{end}}}\int_{\mathbb{R}^{d}}\left[L(t,x,v(t,x))+F(t,x,\rho(t,x))\right]\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}^{d}}G(x,\rho(t_{\text{end}},x))\rho(t_{\text{end}},x)\,\mathrm{d}x\,,$ (19)
@@
Let $z_{0}=x_{0}$, under the vector field $f$, we obtain the probability flow (Chen et al., 2024) of the stochastic dynamic $x_{t}$, given by $\partial_{t}z_{t}=f(t,z_{t})$, which coincides with equation 3a. This deterministic dynamic characterizes the probability distribution $\rho(t,x)$. If $z_{0}\sim\rho_{0}$, then the probability distribution for $z_{t}$ is exactly $\rho(t,\cdot)$. With this transformation, the velocity $v(t,x)$ becomes $f(t,x)+\gamma\nabla_{x}\log\rho(t,x)$, and the MFC problem follows

### v2

@@ Abstract
Classical neural ordinary differential equations (ODEs) are powerful tools for approximating the log-density functions in high-dimensional spaces along trajectories, where neural networks parameterize the velocity fields. We specify a system of neural differential equations representing first- and second-order score functions along trajectories based on deep neural networks. We reformulate the mean field control (MFC) problem with individual noises into an unconstrained optimization problem framed by the proposed neural ODE system. Additionally, we introduce a novel regularization term to enforce characteristics of viscous Hamilton–Jacobi–Bellman (HJB) equations to be satisfied based on the evolution of the second-order score function. Examples include regularized Wasserstein proximal operators (RWPOs), probability flow matching of Fokker–Planck (FP) equations, and linear quadratic (LQ) MFC problems, which demonstrate the effectiveness and accuracy of the proposed method.
keywords:
score function, neural ODE, normalizing flow, mean field control
[…]
[labelUCLA]organization=Department of Mathematics,addressline=University of California, city=Los Angeles, postcode=90095, state=CA, country=USA
[…]
[labelUSC]organization=Department of Mathematics, addressline=University of South Carolina, city=Columbia, postcode=29208, state=SC, country=USA
@@ 1 Introduction
Score functions have been widely used in modern machine learning algorithms, particularly generative models through time-reversible diffusion [1]. The score function can be viewed as a deterministic representation of diffusion in stochastic trajectories [2]. In this representation, one reformulates the Brownian motion by the gradient of the logarithm of the density function, after which deterministic trajectories involving score functions can approximate the probability density function. These properties have inspired algorithms for simulating stochastic trajectories or sampling problems that converge to target distributions [3, 4]. Typical applications include modeling the time evolution of probability densities for stochastic dynamics and solving control problems constrained by such dynamics.
While score functions provide powerful tools for modeling stochastic trajectories, their computations are often inefficient, especially in high-dimensional spaces. Classical methods, such as kernel density estimation (KDE) [5], tend to perform poorly in such settings due to the curse of dimensionality [6].
@@
Related work. The normalizing flow has emerged as a powerful technique for solving inference problems, allowing for the construction of complex probability distributions in a tractable manner. This approach was popularized by works such as [7, 8]. The advent of Neural ODEs [9] has attracted considerable attention in the community, opening up a new paradigm for continuous-time deep learning models. Following this breakthrough, numerous extensions have been proposed to enhance the expressiveness and flexibility of these models, such as the augmented Neural ODEs in [10].
@@
Regarding score-based models, [1] studies the score-based reversible time diffusion models in generative modeling. In this case, the score function often comes from the OU process. One needs to solve the score-matching problem to learn the time-dependent score function. In addition, the first-order score dynamic from neural ODEs has been introduced in [11] to work on the simulation of the Fokker-Planck (FP) equation. Rather than using score dynamics directly to compute the probability flow ODEs associated with FP equations, they employed score-matching methods to achieve efficient computations.
@@
The study of MFC problems has become crucial in the last decade [12, 13, 14]. MFC studies strategic decision-making in large populations where individual players interact through specific mean-field quantities. This control formulation is also useful in generative models [15]. For example, neural network-based methods have been employed to solve MFC problems. [16] provided a comprehensive review of neural network approaches for MFC. For first-order MFC problems, [17, 18] developed Neural ODE-based methods. In parallel, [17] introduced numerical algorithms based on the characteristic lines of the Hamilton-Jacobi-Bellman (HJB) equation. For second-order MFC problems, neural networks have also been employed. [19] utilized generative adversarial networks (GANs) to approximate minimization systems for MFC problems, where one neural network models the population density, and the other parameterizes the value function. [20] proposed a PDE-based iterative algorithm to tackle MFC problems. Along this line, MFC problems are generalizations of dynamical optimal transport problems, known as Benamou-Brenier’s formulas [21]. A famous example is the Jordan-Kinderlehrer-Otto (JKO) scheme [22], a variational time discretization in approximating the gradient drift FP equation. The one-step iteration of the JKO scheme can be viewed as an MFC problem. In machine learning computations, the neural JKO scheme [23, 24] was introduced to compute the FP equation, and [25] extended this to approximate general nonlinear gradient flows through a generalized deep neural JKO scheme. This scheme approximates a deterministic MFC problem in each time interval. Other works that use machine learning methods to solve MFC problems include [26, 27]. Different from previous results, our work designs first- and second-order score dynamics to compute second-order MFC problems.
@@ 2 Score-based neural ODE in continuous time
Let $z_{t}\in\mathbb{R}^{d}$ be a flow of random variable for $t\geq 0$ with initialization $z_{0}=x$, which is obtained from a pushforward map $z_{t}=T(t,x)$ in Lagrangian coordinates. For simplicity, we denote $T_{t}:=T(t,\cdot)$, so that $T_{0}$ is an identity map. Throughout this paper, we assume that $T(\cdot,\cdot)$ is smooth and $T_{t}$ is invertible for all $t\geq 0$. We denote the probability distribution of $z_{t}$ by $\rho_{t}$ or $\rho(t,\cdot)$, then $\rho_{t}=T_{t\#}\rho_{0}$, where # is the pushforward operator for distributions. By the change of variable formula, $\rho$ satisfies the well known Monge–Ampère (MA) equation [28]
@@
We leave the proofs for this equation and all propositions afterwards in A.
@@
$\partial_{t}H_{t}$ $=-\sum_{i=1}^{d}s_{it}\nabla_{z}^{2}f_{i}(t,z_{t})-\nabla_{z}^{2}(\nabla_{z}\cdot f(t,z_{t}))$ (3d)
$\quad-H_{t}\nabla_{z}f(t,z_{t})-\nabla_{z}f(t,z_{t})^{\top}H_{t}\,,$
@@
Example: Centered Gaussian distributions. To illustrate the ODE dynamics (3), we consider a concrete example where the random variable follows a centered Gaussian distribution. Let the pushforward map be linear in $x$, i.e. $T(t,x)=T(t)x$, where $T(t):\mathbb{R}_{+}\to\mathbb{R}^{d\times d}$ is a time-dependent matrix. If we define $A(t):\mathbb{R}_{+}\to\mathbb{R}^{d\times d}$ by $\partial_{t}T(t)=A(t)T(t)$, then the vector field of the state dynamics is $f(t,x)=A(t)x$. Now, let the initial state $z_{0}$ follow a centered Gaussian distribution $N(0,\Sigma(0))$ with covariance matrix $\Sigma(0)$. Under this linear map $T(t)$, the state $z_{t}$ remains a Gaussian distribution, with a time-dependent covariance matrix $\Sigma(t)$. In this case, $z_{t}\sim N(0,\Sigma(t))$, and the time evolution of the covariance matrix follows the matrix ODE (with the derivation in A.1)
@@
In this case, $\nabla_{x}f(t,x)=A(t)$ and $\nabla_{x}\cdot f(t,x)=\operatorname{Tr}(A(t))$, and all higher-order spatial derivatives of $f$ vanish. Therefore, the ODE system (3) simplifies to
@@ 3.1 A generalization of normalizing flow
We recall that [9] interpreted the deep residual neural network as an ODE, with each layer representing one step of the forward Euler scheme. Similar ideas have also been proposed in [29]. In this section, we extend these concepts and demonstrate how the first- and second-order score functions can be computed efficiently using the proposed neural ODE system.
We partition the time interval $[0,t_{\text{end}}]$ into $N_{t}$ subintervals with the length $\Delta t=t_{\text{end}}/N_{t}$ and denote the time stamps by $t_{j}=j\,\Delta t$. We apply the forward Euler scheme to discretize the ODE system (3). We parametrize the vector field $f(t,z;\theta)$ as a neural network with parameter $\theta$. Here, $\theta=[w_{0},w_{1},w_{2},b_{1},b_{2}]\in\mathbb{R}^{k+k\times d+d\times k+k+d}$ and
@@
where $\sigma\colon\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is a vector function with elementwise activation functions, and $k$ is the width of the network. This is the typical structure of a neural network with one hidden layer. Then, we discretize the state dynamic $z_{t}$ through
@@ 3.2 Different algorithms for computing the score function
The expression (3c) or (12) provides a way to compute the score function. In this section, we compare several algorithms for computing the score function, and demonstrate the potential efficiency of using formulas (3c) or (12).
One method is to derive the score function from the pushforward map $T(t,x)$ and the MA equation directly. To be more specific, we apply the operator $\nabla_{x}$ and the logarithm on both sides of the MA equation (1). We obtain
@@
As an alternative, we can compute the score $s_{t}$ through the normalizing flow $l_{t}$, i.e., differentiating $l_{t}=\log\rho(t,z_{t})$ w.r.t $z_{t}$. We assume that the width of the neural network (8) is $k=\mathcal{O}(d)$. The score function can be reformulated as
@@
where $\dfrac{\partial z_{t}}{\partial z_{0}}$ and $\dfrac{\partial l_{t}}{\partial z_{0}}$ are obtained from auto-differentiations of deep neural network functions in (15). Using the chain rule, the total computational computational cost for computing one trajectory of the score function is still $\mathcal{O}(N_{t}\,d^{3})$. Both (17) and (18) require computing the inverse Jacobian of $T$ or solving related linear systems, resulting in a cubic cost in the spatial dimension. As mentioned in [9] (section 4), this is the bottleneck of these methods. We leave more detailed discussions of both methods (17) and (18), and two other methods, to C.
In contrast, the high-order normalizing flow (12) only requires discretizing the ODE dynamics (3a) and (3c), with a total cost of $\mathcal{O}(N_{t}\,d^{2})$ for computing one score trajectory $\{s_{t_{j}}\}_{j=0}^{N_{t}}$. As is shown in table 1, our formulation is more efficient compared with (17) and (18), especially in high dimensions. Additionally, we are able to compute the second-order score function through 13.
pushforward map (17) $\mathcal{O}(N_{t}\,d^{3})$
normalizing flow (18) $\mathcal{O}(N_{t}\,d^{3})$
high-order normalizing flow (12) (ours) $\mathcal{O}(N_{t}\,d^{2})$
@@
Example: Gaussian distribution. Our formulation for the second-order score function also has advantages in the example of Gaussian distributions, where $H_{t}=\Sigma(t)^{-1}$ is exactly the inverse of the covariance matrix. Traditionally, estimating $\Sigma(t)^{-1}$ involves several steps. One needs to sample sufficiently many points $z_{0}^{(n)}$, simulate the state dynamic to obtain $z_{t}^{(n)}$, estimate the covariance matrix from these samples, and finally compute the inverse of this estimation. To achieve an error of $\varepsilon$, at least $\mathcal{O}(\varepsilon^{-2})$ samples are required for accurate covariance estimation. Compared to our second-order score dynamic (3d), we only need to pick $\Delta t=\mathcal{O}(\varepsilon)$ and compute $H_{t}$ through (13), with a total cost of $\mathcal{O}(\varepsilon^{-1})$. This results in a more efficient algorithm for estimating score functions in time-dependent Gaussian distributions.
@@ 4 Solving second-order MFC problems
In this section, we apply the score-based normalizing flow to solve the second-order MFC problem. In particular, we demonstrate that the MFC problem can be represented by the optimal control problem of the neural dynamical system in Proposition 1. For readers interested in the connection between MFC problems and generative AI, we refer to the work of [15]. In the context of machine learning, [30] also studies similar problems, in which they compute Wasserstein Lagrangian flows. In this work, we can deal with the diffusion term in MFC problems using proposed neural normalization flows.
@@ 4.1 Formulation of MFC problem
$\inf_{v}\int_{0}^{t_{\text{end}}}$ $\int_{\mathbb{R}^{d}}\left[L(t,x,v(t,x))+F(t,x,\rho(t,x))\right]\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t$ (19)
@@
Let $z_{0}=x_{0}$, under the vector field $f$, we obtain the probability flow [31] of the stochastic dynamic $x_{t}$, given by $\partial_{t}z_{t}=f(t,z_{t})$, which coincides with (3a). This deterministic dynamic characterizes the probability distribution $\rho(t,x)$. If $z_{0}\sim\rho_{0}$, then the probability distribution for $z_{t}$ is exactly $\rho(t,\cdot)$. With this transformation, the velocity $v(t,x)$ becomes $f(t,x)+\gamma\nabla_{x}\log\rho(t,x)$, and the MFC problem follows